OpenStudy (anonymous):

Estimate the circle to the nearest square inch. 1. 88 in.2 2. 62in.2 3. 616in.2 (got 1. 88in.2 ) but i think im wrong

OpenStudy (anonymous):

@gypsy1274

OpenStudy (whpalmer4):

You're to find the area of the circle. \[A = \pi r^2\]Look at the diagram, figure out the radius \(r\), and plug into the formula. \(\pi \approx 3.14\). Your current answer is incorrect.

OpenStudy (anonymous):

@Jesstho.-. @wolfe8 @whpalmer4 is it 2? , i may be looking at the wrong one again, i dont have my glasses, im hal blindd

OpenStudy (anonymous):

@gypsy1274 @kmullis6 @bbcream14

OpenStudy (anonymous):

give me a sec

OpenStudy (anonymous):

@ElvenBlackout

OpenStudy (whpalmer4):

\[\huge\text{The radius of the circle is 14. }\]\[\huge\text{ What is 3.14*14*14=}\]\[\huge\text{answer choices are 88, 62, 616}\]

OpenStudy (anonymous):

3.14 * 14 * 14 = 615.44 @whpalmer4

OpenStudy (whpalmer4):

\[\huge\text{Yes, that is answer #3, 616 in}^2\text{ after rounding to nearest in}^2\]